OpenStudy (anonymous):

Tell me if this is correct,... Solve for x: 7 over 8 minus 1 over x equals 3 over 4 x = 1 x = 2 x = 4 x = 8 I picked C

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 }-\frac{ 1 }{ x}=\frac{ 3 }{ 4}\]

jimthompson5910 (jim_thompson5910):

let's check C to see if it's the correct answer 7/8 - 1/x = 3/4 7/8 - 1/4 = 3/4 .. plug in x = 4 7/8 - 2/8 = 3/4 (7 - 2)/8 = 3/4 5/8 = 3/4 So because the last equation is false, this means that x = 4 is NOT the solution So C is NOT the answer

OpenStudy (anonymous):

solve the equation and u l get the answer c is not the answer @dreamer05

OpenStudy (luigi0210):

Sorry about that. But yea get rid of the fractions. So pick a common denominator

OpenStudy (whpalmer4):

What is 3/4 if converted to 8ths?

OpenStudy (whpalmer4):

(multiply both top and bottom by 2)

Directrix (directrix):

@dreamer05